OpenStudy (anonymous):

Do anbody kw what 3x+2=7 is?

OpenStudy (uri):

Yes :P

OpenStudy (uri):

3x+2=7 3x =7-2 3x =5 x=5/3 divide and answer :P

OpenStudy (anonymous):

Thnk you uri.

OpenStudy (uri):

yw. ;)